2(x+2(-3/2x-1))= x/3
Found 2 solutions by helpnalgebra, Alan3354:
Answer by helpnalgebra(91)

this is step by step
2(x+2(-3x/2-1)))=x/3
2(x-3x-2)=x/3
2(-2x-2)=x/3
-4x-4=x/3
-4x-3-4•3=x/3•3
-12x-12=x/3•3
-12x-12=x
-12x-12-x=0
-13x-12=0
-13x=12
-13x/-13=12/-13
x=12/-13
x=-12/13

Answer by Alan3354(69443)

2(x+2(-3/2x-1))= x/3
6(x+2(-3/2x-1))= x
6(x + (-3x-2)) = x
6(-2x-2) = x
-12x - 12 = x
13x = -12
x = -12/13